OpenStudy (jordanloveangel):

Which equation shows the substitution method being used to solve the system of linear equations? x + y = 6 x = y + 5 A. x = (x – 6) + 5 B. (y + 5) + y = 6 C. x + (y + 5) = 6 D. x + y = y + 5

OpenStudy (jordanloveangel):

@Godlovesme

OpenStudy (godlovesme):

give me a sec brb @jordanloveangel

OpenStudy (jordanloveangel):

k

OpenStudy (bradely):

x + y = 6 ----- (1) x = y + 5 ----- (2) Substitute x = y + 5 in equation (1) (y + 5) + y = 6 Option B is correct Source: http://www.mathskey.com/

OpenStudy (precal):

when you use the sub method you always have to define one of the variables. Ie x= something or y=something

OpenStudy (precal):

notice the second equation is x= something you will need to sub that equation into the first equation

OpenStudy (precal):

x=y+5 sub y+5 into the x of the first equation. Source: My brain and previous knowledge of how to solve systems